DX LUCI/AO

00:30
  • Setting up for OSCO
1:08
  • Having problems with the preset.
  • Something was stuck and the preset wouldn't go through.
  • Then we didn't find the star. It is 1' off of center, as seen in LUCI N3.75
  • Ilocater found the star and wanted to take a data set, then we'll try again.
2:21 03:42
  • GOPT now running.
  • New preset, mag 5.8
  • 034307, 034601, 034944, 035013
4:05
  • Going to head to hr 8799
04:08
  • Shell rip--Need to restart the AdSec
04:16
  • Back to the new preset
  • 043134, 043217, 043300, 043411
05:17
  • Going to a slightly fainter star
  • HIP 9452
  • 053518, 053543, 053601, 053617
  • Seeing Bubble caused the loop to open.
06:05
  • To HIP 7981
  • 061713, 061735, 061757, 061811

-- GregoryTaylor - 13 Nov 2019
Topic revision: r4 - 13 Nov 2019, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback